OpenStudy (anonymous):

Giving medals!!! Find all solutions to the equation. 7 sin2x - 14 sin x + 2 = -5

OpenStudy (mathmate):

\(7sin^2(x)-14sin(x)+2=-5\) ?

OpenStudy (anonymous):

yeah

OpenStudy (mathmate):

First, we like to transpose -5 to the left: \(7sin^2(x)-14sin(x)+7=0\) Then factor out the value 7, and hence simplify the equation: \(sin^2(x)-2sin(x)+1=0\) Substitute u=sin(x) to reduce the problem to a quadratic: \(u^2-2u+1=0\) Factorize expression on the left: \((u-1)^2=0\) that gives a solution of u=1 which means sin(x)=1. Can you solve this equation (completely) now?

OpenStudy (anonymous):

yes thats exactly what I needed! thanks!

OpenStudy (mathmate):

You're welcome! :)